Solve for debt and equity rates. Step by step please!

7.2% = We(5.28%) + We(10%)
Answer #1

Let the weight of equity be"X" so weight of Debt = 1-X

7.2% = (1-X) *5.28 % + (X *10%)

7.2 = 5.28 - 5.28X + 10X

7.2 -5.28 = 4.72 X

1.92 /4.72 = X

X= .4068 or 40.68%

Weight of equity = .4068 or 40.68%

Weight of debt = 1- .4068   = .5932 or 59.32%